Agenda/Minutes of the VIS Meeting on 2020/8/7
2020/8/7 14:00 - 14:50
Participants: Takahashi, Fabian, Miyo, Aso, Hirata, Terrence, Shoda, Sato, Tanaka, Washimi
[Miyo] by 08/28/2020?, 20% done (+0%), Make a cabling diagram for O4
[Miyo] by 08/28/2020?, 25% done (+0%), Make prototype system for temperature control
[Miyo] by 08/28/2020?, 20% done (+0%), Enable to control the limit switch of stepper motor
[Miyo] by 08/28/2020?, 0% done (+0%), Request AEL chassis for picomotor driver
[Miyo] by 08/28/2020?, 0% done (+0%), Update remote control system
[Miyo] by 08/28/2020?, 0% done (+0%), Add CRY stepper motor in the remote control system
- (This week) Installation was finished. Now I try to diagonalize sensors and balance actuators by using the Prequa system.
- (Next week) I will do the modal damping on TypeA suspensions after the initialization work for sensors and actuators.
On Monday I submitted the paper to CPC: JGW-P2011860.
Noise budget paper: I calculated upper limits to DAC and coil driver noise contribution to DARM for the BS: klog 14875.
Noise budget paper: Fabian calculated upper limits to DAC and coil driver noise contribution to DARM for PRM: klog 14876.
- Brainstorming meeting for IMMT2 modification was held yesterday. We discussed the modified shape of the outer frame and how to replace it. The plan is to replace the frame without removing IMMT2 from the chamber.
- Updated the sensor correction path in the Type-A modeling. The simulated signals at IP stage fitted well to the actual signal when the sensor correction accuracy is about 90 %.
With the updated control model, the TM displacement at micro-seismic motion was reduced by about 10 times. Therefore, the actuation efficiency should be about 10 times smaller than the design value to be consistent with the measured FB signal. If I roughly estimate the actuation efficiency, it is almost consistent with what CAL team measured.14864
I observed the pendulum wobbling when the asymmetry is introduced in the suspension wires. 14868
- Evaluated the modified LVDT cards (grounding of the demodulator chip).
- KAGRA is in "Upgrade phase for O4". Prequa system was implemented to the Type-A.
- Confirmation of personnel at the site.
- Type-A team: Takahashi, Sato, Miyo, + Company(Jan-Apr)
- Type-B/Bp team: (Shoda,) Hirata, Fabian, Washimi, + Company(Oct-Dec)
- Procedure documents are required by SEO.
Schedule towards O4 (Chart) was revised.
[Washimi] Room Temmperature report(https://www.icrr.u-tokyo.ac.jp/~washimi/KAGRA/SEM/fig_GAS/)
- On 2020/8/19(Wed)